OpenStudy (anonymous):

8=8p+13-3p P=? What does P equal to?

OpenStudy (anonymous):

Subtract 8p-3p = 5p 8 = 13 + 5p Subtract 13 from both sides 8 - 13 = 13 - 13 + 5p -5 = 5p -1 = p

OpenStudy (anonymous):

Thankyou very much!

OpenStudy (anonymous):

:)